OpenStudy (anonymous):

I am having trouble converting with bases. Can anyone help me with this and explain it better then the book. Convert a3 fourteen to a base six.

OpenStudy (rational):

convert to base10 first

OpenStudy (rational):

\[\large \left(A3\right)_{14} = 10\times 14^1 + 3\times 14^0 = \left(143\right)_{10}\]

OpenStudy (rational):

convert base10 to base6

OpenStudy (anonymous):

ok - I see how you converted to the base 10 and I understand that. Our book is not very good and does not give many examples.

OpenStudy (rational):

good :) do you know how to convert it to base6 ?

OpenStudy (anonymous):

I do not...but I am re-reading now on how to convert a base 10 to another base

OpenStudy (rational):

its easy, you just need to count in powers of 6 : ... 6^3, 6^2, 6^1, 6^0

OpenStudy (rational):

how many 6^3's are there in 143 ?

OpenStudy (anonymous):

ok - I am following

OpenStudy (rational):

in other words, how many times 6^3 goes in 143 ?

OpenStudy (rational):

there are 0 6^3's in 143, right ?

OpenStudy (anonymous):

would it be 78six?

OpenStudy (rational):

it has to occupy more positions in base6

OpenStudy (rational):

i mean it has to occupy atleast 3 digits in base6

OpenStudy (anonymous):

ok - let me retry

OpenStudy (anonymous):

103 six?

OpenStudy (rational):

convert it to base10, do u get 143 back ?

OpenStudy (rational):